How do you solve the following equation using the Elimination Method:

-3x + 3y = 3
3x + y = 9

Answer by robertb(5830)

Add corresponding sides of the two equations. -3x and 3x will cancel each other out, giving 3y+y = 3+9, or 4y = 12, or y = 3. Using the 2nd equation, 3x + 3 = 9, or 3x = 6, or x = 2. Therefore the solution is the ordered pair (2, 3).